# Very-long-baseline interferometry study of the flaring blazar TXS 1508+572 in the early Universe

### v2

@@ Abstract
Results. We present total intensity images from our multi-wavelength VLBI monitoring that reveal significant morphological changes in the parsec-scale structure of TXS 1508+572. The jet proper motion values range from $0.12$ mas/yr to $0.27$ mas/yr, which corresponds to apparent superluminal motion $\beta_{\mathrm{app}}\approx 14.3\,c$$32.2\,c$. This is consistent with the high Lorentz factors inferred from the spectral energy distribution (SED) modeling for this source. The core shift measurement reveals no significant impact by the high-energy flare on the distance of the $43$-GHz radio core with respect to the central engine, that means this region is probably not affected by e.g., injection of new plasma as seen in other well-studied sources like CTA 102. We determine the average distance from the $43$-GHz radio core to the central supermassive black hole to be $46.1\pm 2.3\leavevmode\nobreak\ \mu$as, that corresponds to a projected distance of $0.32\pm 0.02$ pc. We estimate the equipartition magnetic field strength $1$ pc from the central engine to be on the order of $0.8$ G, and the non-equipartition magnetic field strength at the same distance to be about $1.7$ G, which values agree well with the magnetic field strength measured in low to intermediate redshift AGN.
@@ 3.2 Spectral index maps and core shift measurements
The equipartition magnetic field strengths, $B^{\mathrm{eq}}_{\mathrm{1pc}}$, derived for the three considered epochs are the following: $0.86\pm 0.05$ G, $0.71\pm 0.08$ G, and $0.81\pm 0.07$ G. The $B^{\mathrm{non-eq}}_{\mathrm{1pc}}$ values for the same epochs are $1.7\pm 0.8$ G, $0.6\pm 0.4$ G, and $1.8\pm 1.2$ G, respectively.
In order to compare $B^{\mathrm{eq}}$ to that of M 87 on horizon scales (Event Horizon Telescope Collaboration et al. 2021), we extrapolate the average $B^{\mathrm{eq,aver}}=0.79\pm 0.04$ G to $5\leavevmode\nobreak\ r_{\mathrm{g}}$ or $0.0036$ pc projected distance. The gravitational radius is calculated as $r_{\mathrm{g}}=GM_{\mathrm{BH}}/c^{2}$, where $G$ is the gravitational constant and $M_{\mathrm{BH}}$ is the black hole mass assumed to be $1.5\times 10^{10}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ based on the SED fit parameters in Gokus et al. (2024). We measure $B^{\mathrm{eq}}_{\mathrm{5r_{g}}}=220.4\pm 10.9$ G, that is significantly larger than the $B\sim 1-30$ G reported for M 87 (Event Horizon Telescope Collaboration et al. 2021).
@@ 4.3 Core shift evolution
Based on the core shift measurement, we derived the equipartition magnetic field strengths $1$ pc from the SMBH (Zdziarski et al. 2015), which are $0.86\pm 0.05$ G for the first, $0.71\pm 0.08$ G for the second, and $0.81\pm 0.07$ G for the third epoch. These magnetic field strengths are close to the ones derived for the MOJAVE sample (Pushkarev et al. 2012) and by Zamaninasab et al. (2014) for AGN below redshift $2.43$ and $2.37$, respectively. Non-equipartition magnetic field strengths $1$ pc from the SMBH are $1.7\pm 0.8$ G, $0.6\pm 0.4$ G, and $1.8\pm 1.2$ G for the first, second, and third epochs, respectively. Extrapolating the average $B^{\mathrm{eq}}_{\mathrm{1pc}}$ to $5r_{\mathrm{g}}$, we measure $B^{\mathrm{eq}}_{\mathrm{5r_{g}}}=220.4\pm 10.9$ G, a value significantly larger than the $B\sim 1-30$ G reported for M 87 on horizon scales (Event Horizon Telescope Collaboration et al. 2021).
@@ 5 Summary
Using our multi-frequency data, we measured the distance to the central engine based on the core shift. The distance to the jet apex stays consistent within the measurement errors throughout our observations. On average, the central engine is located $46.1\pm 2.3\leavevmode\nobreak\ \mu$as or $0.32\pm 0.02$ pc from the $43$-GHz VLBI core. Under the equipartition assumption, which is supported by our brightness temperature measurements in the presence of Doppler boosting, we measure $B^{\mathrm{eq}}_{\mathrm{1pc}}$ of $0.86\pm 0.05$ G, $0.71\pm 0.08$ G, and $0.81\pm 0.07$ G for the three epochs. Low to intermediate redshift AGN also exhibit similar values of $B^{\mathrm{eq}}_{\mathrm{1pc}}$ (Pushkarev et al. 2012; Zamaninasab et al. 2014). $B^{\mathrm{non-eq}}_{\mathrm{1pc}}$ values are somewhat higher than this, with $1.7\pm 0.8$ G, $0.6\pm 0.4$ G, and $1.8\pm 1.2$ G for the consecutive epochs.

### v3

@@ Abstract
Results. We present total intensity images from our multi-wavelength VLBI monitoring that reveal significant morphological changes in the parsec-scale structure of TXS 1508+572. The jet proper motion values range from $0.12$ mas/yr to $0.27$ mas/yr, which corresponds to apparent superluminal motion $\beta_{\mathrm{app}}\approx 14.3\,c$$32.2\,c$. This is consistent with the high Lorentz factors inferred from the spectral energy distribution (SED) modeling for this source. The core shift measurement reveals no significant impact by the high-energy flare on the distance of the $43$-GHz radio core with respect to the central engine, that means this region is probably not affected by e.g., injection of new plasma as seen in other well-studied sources like CTA 102. We determine the average distance from the $43$-GHz radio core to the central supermassive black hole to be $46.1\pm 2.3\leavevmode\nobreak\ \mu$as, that corresponds to a projected distance of $0.32\pm 0.02$ pc. We estimate the equipartition magnetic field strength $1$ pc from the central engine to be on the order of $1.8$ G, and the non-equipartition magnetic field strength at the same distance to be about $257$ G, the former of which values agrees well with the magnetic field strength measured in low to intermediate redshift AGN.
@@ 3.2 Spectral index maps and core shift measurements
The equipartition magnetic field strengths, $B^{\mathrm{eq}}_{\mathrm{1pc}}$, derived for the three considered epochs are the following: $1.91\pm 0.11$ G, $1.57\pm 0.17$ G, and $1.79\pm 0.16$ G. The $B^{\mathrm{non-eq}}_{\mathrm{1pc}}$ values for the same epochs are $324.7\pm 156.3$ G, $104.9\pm 82.7$ G, and $341.9\pm 225.2$ G, respectively.
In order to compare $B^{\mathrm{eq}}$ to that of M 87 on horizon scales (Event Horizon Telescope Collaboration et al. 2021), we extrapolate the average $B^{\mathrm{eq,aver}}=0.79\pm 0.04$ G to $5\leavevmode\nobreak\ r_{\mathrm{g}}$ or $0.0036$ pc projected distance. The gravitational radius is calculated as $r_{\mathrm{g}}=GM_{\mathrm{BH}}/c^{2}$, where $G$ is the gravitational constant and $M_{\mathrm{BH}}$ is the black hole mass assumed to be $1.5\times 10^{10}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ based on the SED fit parameters in Gokus et al. (2024). We measure $B^{\mathrm{eq}}_{\mathrm{5r_{g}}}=487.4\pm 33.0$ G, that is significantly larger than the $B\sim 1-30$ G reported for M 87 (Event Horizon Telescope Collaboration et al. 2021).
@@ 4.3 Core shift evolution
Based on the core shift measurement, we derived the equipartition magnetic field strengths $1$ pc from the SMBH (Zdziarski et al. 2015), which are $1.91\pm 0.11$ G for the first, $1.57\pm 0.17$ G for the second, and $1.79\pm 0.16$ G for the third epoch. These magnetic field strengths are close to the ones derived for the MOJAVE sample (Pushkarev et al. 2012) and by Zamaninasab et al. (2014) for AGN below redshift $2.43$ and $2.37$, respectively. Non-equipartition magnetic field strengths $1$ pc from the SMBH are $324.7\pm 156.3$ G, $104.9\pm 82.7$ G, and $341.9\pm 225.2$ G for the first, second, and third epochs, respectively. Extrapolating the average $B^{\mathrm{eq}}_{\mathrm{1pc}}$ to $5r_{\mathrm{g}}$, we measure $B^{\mathrm{eq}}_{\mathrm{5r_{g}}}=487.4\pm 33.0$ G, a value significantly larger than the $B\sim 1-30$ G reported for M 87 on horizon scales (Event Horizon Telescope Collaboration et al. 2021).
@@ 5 Summary
Using our multi-frequency data, we measured the distance to the central engine based on the core shift. The distance to the jet apex stays consistent within the measurement errors throughout our observations. On average, the central engine is located $46.1\pm 2.3\leavevmode\nobreak\ \mu$as or $0.32\pm 0.02$ pc from the $43$-GHz VLBI core. Under the equipartition assumption, which is supported by our brightness temperature measurements in the presence of Doppler boosting, we measure $B^{\mathrm{eq}}_{\mathrm{1pc}}$ of $1.91\pm 0.11$ G, $1.57\pm 0.17$ G, and $1.79\pm 0.16$ G for the three epochs. Low to intermediate redshift AGN also exhibit similar values of $B^{\mathrm{eq}}_{\mathrm{1pc}}$ (Pushkarev et al. 2012; Zamaninasab et al. 2014). $B^{\mathrm{non-eq}}_{\mathrm{1pc}}$ values are significantly higher than this, with $324.7\pm 156.3$ G, $104.9\pm 82.7$ G, and $341.9\pm 225.2$ G for the consecutive epochs.